OpenStudy (anonymous):

Given 8r+3s=12 and 7r+2s=9, find the value of 5(r+s)

OpenStudy (anonymous):

a)5 b)10 c)15 d)20 e)25

OpenStudy (jazzyfa30):

@Trai7

OpenStudy (anonymous):

subtract second equation from first equation and multiply by 5

OpenStudy (anonymous):

8r+3s-7r-2s=12-9 r+s=3

OpenStudy (anonymous):

you 15 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanx ^^

OpenStudy (anonymous):

yw